# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ IV-B3 One-dimensional Parameter Estimation Algorithms
The sequential domain method first performs AoA estimation once, followed by performing spatial domain beamforming and delay estimation $K_{\hat{\theta}}$ times, and finally delay domain beamforming and Doppler estimation $K$ times. Compared to sequential one-domain estimation, parallel one-domain estimation requires performing AoA, delay, and Doppler estimation only once, thus reducing complexity. However, the parallel method requires the extra step of parameter grouping, whereas the sequential method associates the parameters automatically. If the sensing targets cannot be distinguished in a single domain, the parallel method may result in poor estimation performance caused by a failure in parameter grouping. In contrast, the sequential method can use beamforming to extract the signal of each target as long as the targets can be distinguished based on one parameter, thus enabling accurate estimation of the other two target parameters. The sequential method can also effectively prevent weaker targets from being obscured by the sidelobes of stronger targets by first separating out the signals of the stronger targets [113]. Therefore, the parallel method is preferred if the AoA, delay, and Doppler resolutions are all high; otherwise, the sequential method is preferable. Moreover, for sequential one-domain parameter estimation, since the failure to distinguish different targets before performing beamforming will degrade estimation performance, the target parameter with the highest resolution should be estimated first, followed by that with the second highest one and finally that with the lowest one.
@@ IV-C1 Joint Two-domain Parameter Estimation Model
In this subsection, we present the joint two-domain parameter estimation methods shown in Fig. 4. Compared to the decoupled parameter estimation discussed in Section IV-B, joint two-domain parameter estimation can achieve higher accuracy at the cost of higher computational complexity. The two-dimensional (2D) joint target parameter estimation can be divided into three cases, corresponding to joint delay-Doppler estimation, joint AoA-delay estimation, and joint AoA-Doppler estimation. Since the three cases have mathematically similar forms, we focus on the joint delay-Doppler estimation, while all signal processing techniques and algorithms can be similarly applied to joint AoA-delay estimation and joint AoA-Doppler estimation.
As illustrated in Fig. 4, before performing joint delay-Doppler estimation, we first estimate the AoAs of the sensing targets based on (22), and then perform beamforming on (LABEL:d25) to decouple signals from different angles. However, joint two-domain parameter estimation aims to jointly estimate the delay and Doppler for the targets at each angle, rather than sequentially estimating these two parameters. Assume that $K_{\hat{\theta}}$ AoAs are estimated, with $N_{k_{\hat{\theta}}}$ targets at angle $\hat{\theta}_{k_{\hat{\theta}}}$. After spatial domain beamforming, the extracted signals $\mathbf{X}_{k_{\hat{\theta}}}$ in (38) for each estimated angle are exploited to simultaneously estimate delay $\tau_{k}$ and Doppler frequency shift $\upsilon_{k}$ of the sensing targets.
@@
2D-Periodogram: For the joint delay-Doppler estimation problem, based on the signal in (38), the 2D power spectrum $f(\tau,\upsilon)$ in the delay-Doppler domain with noise ignored can be obtained in (63) as shown on top of the next page,
@@
where $\mathbf{a}_{\tau}(\tau)$ and $\mathbf{a}_{\upsilon}(\upsilon)$ are the steering vectors in the delay and Doppler domains in (27) and (31), $\tau\in\left[0,\frac{1}{\Delta f}\right]$ and $\upsilon\in\left[-\frac{1}{2T_{s}},\frac{1}{2T_{s}}\right]$ are the observation delay and Doppler frequency, respectively. The approximation $(a)$ in (63) holds if the delay or Doppler difference between sensing targets exceeds the corresponding resolution. (63) reveals that the power spectrum in the delay-Doppler domain is the product of the powers in the two domains, and the two components are mutually independent. Therefore, as long as the targets can be distinguished in either domain, the power spectrum can be used to distinguish the targets.
Similar to the 1D-periodogram, to reduce the computational complexity of searching $f(\tau,\upsilon)$ over observation delay $\tau$ and Doppler $\upsilon$, an $N_{\tau}^{\mathrm{IFFT}}$-point IFFT and $N_{\upsilon}^{\mathrm{FFT}}$-point FFT can be performed in the subcarrier and symbol domains of $\mathbf{X}_{k_{\hat{\theta}}}$ in (38) for fast computation. When ${N_{\tau}^{\mathrm{IFFT}}}>N$ or ${N_{\upsilon}^{\mathrm{FFT}}}>P$, zero-padding is utilized to convert $\mathbf{X}_{k_{\hat{\theta}}}\in\mathbb{C}^{N\times P}$ to $\mathbf{X}_{k_{\hat{\theta}}}^{\mathrm{ZP}}\in\mathbb{C}^{{N_{\tau}^{\mathrm{IFFT}}}\times{N_{\upsilon}^{\mathrm{FFT}}}}$. Then, the resulting 2D-periodogram $\mathbf{F}\in\mathbb{C}^{{N_{\tau}^{\mathrm{IFFT}}}\times{N_{\upsilon}^{\mathrm{FFT}}}}$ can be expressed as
@@
where $\mathbf{W}^{*}$ is the conjugate of IDFT matrix $\mathbf{W}$. Negative Doppler frequency estimation necessitates a spectral shift in the Doppler domain of $\mathbf{F}$ to center the zero frequency component within the spectrum. Then, the peak index $\big{(}\hat{n}_{\tau}^{i,{k_{\hat{\theta}}}},\hat{n}_{\upsilon}^{i,{k_{\hat{\theta}}}}\big{)}$ of $\mathbf{F}$ indicates the phases $\Delta f\hat{\tau}_{i,{k_{\hat{\theta}}}}=\frac{\hat{n}^{i,{k_{\hat{\theta}}}}_{\tau}}{{N_{\tau}^{\mathrm{IFFT}}}}$ and $T_{s}\hat{\upsilon}_{i,{k_{\hat{\theta}}}}=\frac{\hat{n}^{i,{k_{\hat{\theta}}}}_{\upsilon}}{{N_{\upsilon}^{\mathrm{FFT}}}}$ in the delay and Doppler domains, respectively. Thus, the delay $\hat{\tau}_{i,{k_{\hat{\theta}}}}$ and Doppler frequency $\hat{\upsilon}_{i,{k_{\hat{\theta}}}}$, $\forall i=1,\cdots,{N_{k_{\hat{\theta}}}}$, of the $N_{k_{\hat{\theta}}}$ sensing targets at angle $\hat{\theta}_{k_{\hat{\theta}}}$ can be estimated as
@@
2D-MUSIC: In order to apply super-resolution algorithms to the joint delay-Doppler estimation problem, the data matrix $\mathbf{X}_{k_{\hat{\theta}}}\in\mathbb{C}^{N\times P}$ in (38) is vectorized as $\mathbf{x}_{k_{\hat{\theta}}}\in\mathbb{C}^{NP\times 1}$, leading to
$\mathbf{x}_{k_{\hat{\theta}}}$ $=\operatorname{vec}\left(\mathbf{X}_{k_{\hat{\theta}}}\right)$ (66)
$=\sum_{i=1}^{N_{k_{\hat{\theta}}}}\tilde{\alpha}_{i,{k_{\hat{\theta}}}}\mathbf{a}_{\upsilon}(\upsilon_{i,{k_{\hat{\theta}}}})\otimes\mathbf{a}_{\tau}(\tau_{i,{k_{\hat{\theta}}}})+\operatorname{vec}\left(\mathbf{Z}_{{k_{\hat{\theta}}}}\right).$
@@
where $\mathbf{E}_{n}$ denotes the noise subspace, $\mathbf{a}_{S}\left(\tau,\upsilon\right)=\mathbf{a}_{\upsilon}\left(\upsilon,P_{\mathrm{sub}}\right)\otimes\mathbf{a}_{\tau}\left(\tau,N_{\mathrm{sub}}\right)\in\mathbb{C}^{N_{\mathrm{sub}}P_{\mathrm{sub}}\times 1}$ is the corresponding steering vector, $\mathbf{a}_{\tau}\left(\tau,N_{\mathrm{sub}}\right)$ and $\mathbf{a}_{\upsilon}\left(\upsilon,P_{\mathrm{sub}}\right)$ are the steering vectors in the delay and Doppler domains in (27) and (31). The peaks of $P_{\mathrm{MUSIC}}\left(\tau,\upsilon\right)$ correspond to the delays and Doppler frequency shifts of the sensing targets.
(a) 2D-Periodogram spectrum.
(b) 2D-MUSIC spectrum.
For the simulation setup in Table III, consider the challenging scenario where the AoAs of the three targets are all $10^{\circ}$, while the other parameters remain unchanged. A comparison of the 2D-periodogram and 2D-MUSIC for joint delay-Doppler estimation in this case is shown in Figs. 10(a) and 10(b), where MSSP is perfomed for the 2D-MUSIC algorithm with a window size of $N_{\mathrm{sub}}\times P_{\mathrm{sub}}=\frac{N}{2}\times\frac{P}{2}=64\times 32$. Similar to the 1D-periodogram and 1D-MUSIC, the spectrum of 2D-MUSIC is smoother and exhibits sharper peaks compared to that of the 2D-periodogram. Consequently, in principle 2D-MUSIC offers higher delay and Doppler resolution and can prevent weak targets from being obscured by the sidelobes of strong targets, thereby reducing inter-target interference and achieving better sensing accuracy. However, for 2D-MUSIC, one must correctly choose the dimension of the signal subspace in order to include the impact of the weak targets, and this is a difficult task in general.
@@ IV-D Joint Three-domain Parameter Estimation
In this subsection, we present the joint three-domain parameter estimation method shown in Fig. 4. The joint AoA-delay-Doppler estimation simultaneously determines the angles, delays, and Doppler frequency shifts of the sensing targets using the signal $\overline{\mathbf{Y}}_{s}$ after element-wise symbol division in (LABEL:d25).
@@
Similar to 1D and 2D cases, a spectral shift in the angle and Doppler domains of $\mathbf{F}_{\mathrm{3D}}$ to center the zero frequency component within the spectrum is required. Then, from the peak indices $\big{(}\hat{n}^{k}_{\theta},\hat{n}^{k}_{\tau},\hat{n}^{k}_{\upsilon}\big{)}$ of $\mathbf{F}_{\mathrm{3D}}$, the angle $\hat{\theta}_{k}$, delay $\hat{\tau}_{k}$, and Doppler frequency $\hat{\upsilon}_{k}$, $\forall k=1,\cdots,K$, of the $K$ sensing targets can be estimated as
@@
The ESPRIT and OMP algorithms can also be applied to the 2D and 3D parameter estimation problems. Similar to the MUSIC algorithm, spatial smoothing is also required for ESPRIT to achieve decorrelation. The conventional 2D and 3D ESPRIT algorithms require parameter grouping [115] to match the estimation results of the different domains, whereas some modified ESPRIT algorithms [116, 117] can achieve automatic grouping. Furthermore, extensive efforts have been devoted to reducing the complexity for super-resolution methods. For example, the ROOT-MUSIC [118, 119] algorithm replaces the spectrum search in MUSIC with polynomial rooting, the propagator method (PM) [120, 121] uses an alternative to achieve the eigenvalue decomposition, and the FFT-MUSIC [122] algorithm accelerates the MUSIC spectrum peak search using FFT operations. In the following, we analyze the complexity and estimation accuracy of the aforementioned algorithms.
IV-E Complexity and Performance Analysis
IV-E1 Complexity Analysis
We examine the theoretical computational complexity of the aforementioned algorithms in terms of their required number of complex-valued multiplications. As such, the complexity of the multiplication of $M\times Q$ and $Q\times N$ matrices is $\mathcal{O}(QMN)$, and that of the EVD of an $M\times M$ matrix is $\mathcal{O}(M^{3})$. The complexity of various 1D parameter estimation algorithms for AoA estimation is shown in Table IV, while these results also apply to delay and Doppler estimation. In these expressions, $Q$ denotes the number of snapshots, $N^{\mathrm{FFT}}/N^{\mathrm{IFFT}}$ and $N^{\mathrm{s}}$ denote the number of FFT/IFFT points and spectrum search points respectively, which are determined by the estimation interval $\eta$, and $N^{\mathrm{FFT}}=N^{\mathrm{IFFT}}=N^{\mathrm{s}}=\frac{\Delta_{\mathrm{max}}}{\eta}$ with $\Delta_{\mathrm{max}}$ representing the maximum estimation range. The estimation interval affects estimation accuracy, and it should be adjustable in order to achieve specific accuracy requirements. To our best knowledge, it is difficult to express the theoretical complexity of ROOT-MUSIC due to the difficulty of quantifying the complexity of the polynomial rooting process. For a typical example, we set the number of OFDM subcarriers and symbols as $N=128$ and $P=64$, respectively, the number of targets is $K=3$, the estimation interval is $\eta=0.1^{\circ}$, and the estimation range is $[-90^{\circ},90^{\circ}]$. Then, the theoretical complexity of the various 1D estimation algorithms for different numbers of Sen-RX antennas $M_{r}$ is shown in Table V.
TABLE IV: Complexity of 1D parameter estimation algorithms for AoA estimation in MIMO-OFDM ISAC.
@@ IV-E1 Complexity Analysis
$Q{M_{r}^{2}}+M_{r}K\left({M_{r}-K}\right)+$
$\left({2M_{r}\left({M_{r}-K}\right)+M_{r}}\right){N_{\theta}^{\text{s}}}$
@@
$Q{M_{r}^{2}}+M_{r}K\left({M_{r}-K}\right)+$
${Q^{2}}\left({M_{r}-1}\right)+{K^{3}}$
@@
TABLE V: The complexity of 1D algorithms for different numbers of antennas for AoA estimation in MIMO-OFDM ISAC.
Algorithm Basic setup and complexity
$N=128,P=64,Q=NP=8192,$
$K=3,N^{s}=N_{\theta}^{\mathrm{IFFT}}=1800$
$M_{r}=16$ $M_{r}=64$ $M_{r}=256$ $M_{r}=512$
Periodogram 1.596e8 1.596e8 1.596e8 1.596e8
MUSIC 2.879e6 4.799e7 7.874e8 3.221e9
PM-MUSIC 2.876e6 4.774e7 7.708e8 3.088e9
FFT-MUSIC 2.354e6 3.500e7 5.586e8 2.292e9
LS-ESPRIT 2.101e6 3.382e7 5.537e8 2.282e9
TLS-ESPRIT 2.102e6 3.382e7 5.537e8 2.282e9
PM-ESPRIT 2.098e6 3.357e7 5.371e8 2.148e9
OMP 7.094e8 2.837e9 1.135e10 2.270e10
As shown in Table V, for MIMO-OFDM ISAC systems, when the numbers of antennas is small, while the number of subcarriers and OFDM symbols are large, numerous snapshots are provided for AoA estimation, and the periodogram algorithm even exhibits higher complexity than MUSIC and ESPRIT. However, the situation changes as the number of antennas increases, the periodogram becomes more computationally efficient due to the fast computation of the FFT, followed by ESPRIT and MUSIC, while OMP has the highest complexity. The periodogram and OMP algorithms can estimate the desired parameters using only partial snapshots or even a single snapshot to reduce complexity, at the cost of increased error. Among MUSIC’s variants, FFT-MUSIC has the lowest complexity, followed by ROOT-MUSIC, with PM-MUSIC offering the least reduction in complexity. Total LS (TLS)-ESPRIT [80] is a variation of the LS-ESPRIT algorithm that offers slightly improved performance, but with a corresponding slight increase in computation. Unlike MUSIC, the PM algorithm has the potential to significantly reduce the complexity of ESPRIT. Overall, the periodogram, FFT-MUSIC, and PM-ESPRIT algorithms offer the lowest complexity.
The complexity analysis for the 2D/3D algorithms follows a similar approach as 1D algorithms and will not be elaborated further here. Note that 2D/3D algorithms not only have a high computational complexity but also involve calculations with large-dimensional matrices, imposing significant demands on computer hardware. To achieve high-accuracy parameter estimation while mitigating these challenges, coarse-fine grid search [125] and signal decimation [81, 126] techniques can significantly reduce both computational complexity and hardware requirements.
(a) RMSE of AoA estimation.
(b) RMSE of delay estimation.
(c) RMSE of Doppler estimation.

### v2

@@ IV-B3 One-dimensional Parameter Estimation Algorithms
The sequential domain method first performs AoA estimation once, followed by spatial domain beamforming and delay estimation $K_{\hat{\theta}}$ times, and finally delay domain beamforming and Doppler estimation $K$ times. Compared to sequential one-domain estimation, parallel one-domain estimation requires performing AoA, delay, and Doppler estimation only once, thus reducing complexity. However, the parallel method requires the extra step of parameter grouping, whereas the sequential method associates the parameters automatically. If the sensing targets cannot be distinguished in a single domain, the parallel method may result in poor estimation performance. In contrast, the sequential method can use beamforming to extract the signal of each target as long as they can be distinguished based on a single parameter, thus enabling accurate estimation of the other two target parameters. The sequential method can also effectively prevent weaker targets from being obscured by first separating out the signals of the stronger targets [124]. Therefore, the parallel method is preferred if the AoA, delay, and Doppler resolutions are all high; otherwise, the sequential method is preferable. Moreover, for sequential one-domain parameter estimation, since the failure to distinguish different targets before performing beamforming will degrade estimation performance, the target parameter with the highest resolution should be estimated first, followed by the parameter with the next highest resolution, and so on.
@@ IV-C1 Joint Two-domain Parameter Estimation Model
In this subsection, we present the joint two-domain parameter estimation methods outlined in Fig. 6. Compared to the decoupled parameter estimation discussed in Section IV-B, joint two-domain parameter estimation can achieve higher accuracy at the cost of higher computational complexity. The two-dimensional (2D) joint target parameter estimation can be divided into three cases, corresponding to joint delay-Doppler estimation, joint AoA-delay estimation, and joint AoA-Doppler estimation. Since the three cases have similar mathematical forms, we will focus only on the joint delay-Doppler estimation, while the other two cases can be handled using similar methods.
As illustrated in Fig. 6, before performing joint delay-Doppler estimation, we first estimate the AoAs of the sensing targets based on (22), and then perform beamforming on (LABEL:d25) to decouple signals from different angles. However, joint two-domain parameter estimation aims to jointly estimate the delay and Doppler for the targets at each angle, rather than sequentially estimating these two parameters. Assume that $K_{\hat{\theta}}$ AoAs are estimated, with $N_{k_{\hat{\theta}}}$ targets at angle $\hat{\theta}_{k_{\hat{\theta}}}$. After spatial domain beamforming, the extracted signals $\mathbf{X}_{k_{\hat{\theta}}}$ in (38) for each estimated angle are exploited to simultaneously estimate the target delay $\tau_{k}$ and Doppler frequency shift $\upsilon_{k}$.
@@
2D-Periodogram: Based on the signal in (38), the noise-free 2D power spectrum $f(\tau,\upsilon)$ in the delay-Doppler domain can be obtained in (63) as shown at the top of the next page,
@@
where $\mathbf{a}_{\tau}(\tau)$ and $\mathbf{a}_{\upsilon}(\upsilon)$ are the steering vectors in the delay and Doppler domains in (27) and (31), $\tau\in\left[0,\frac{1}{\Delta f}\right]$ and $\upsilon\in\left[-\frac{1}{2T_{s}},\frac{1}{2T_{s}}\right]$ are the observation delay and Doppler frequency, respectively. The approximation $(a)$ in (63) holds if the delay or Doppler difference between sensing targets exceeds the corresponding resolution. Eq. (63) reveals that the power spectrum in the delay-Doppler domain is the product of the powers in the two domains, and the two components are mutually independent. Therefore, as long as the targets can be distinguished in either domain, the power spectrum can be used to distinguish them.
Similar to the 1D-periodogram, to reduce the computational complexity of searching $f(\tau,\upsilon)$ over the delay $\tau$ and Doppler $\upsilon$, an $N_{\tau}^{\mathrm{IFFT}}$-point IFFT and $N_{\upsilon}^{\mathrm{FFT}}$-point FFT can be performed in the subcarrier and symbol domains of $\mathbf{X}_{k_{\hat{\theta}}}$ in (38) for fast computation. When ${N_{\tau}^{\mathrm{IFFT}}}>N$ or ${N_{\upsilon}^{\mathrm{FFT}}}>P$, zero-padding is used to convert $\mathbf{X}_{k_{\hat{\theta}}}\in\mathbb{C}^{N\times P}$ to $\mathbf{X}_{k_{\hat{\theta}}}^{\mathrm{ZP}}\in\mathbb{C}^{{N_{\tau}^{\mathrm{IFFT}}}\times{N_{\upsilon}^{\mathrm{FFT}}}}$. Then, the resulting 2D-periodogram $\mathbf{F}\in\mathbb{C}^{{N_{\tau}^{\mathrm{IFFT}}}\times{N_{\upsilon}^{\mathrm{FFT}}}}$ can be expressed as
@@
[…]
@@
(a) 2D-Periodogram spectrum.
(b) 2D-MUSIC spectrum.
2D-MUSIC: To apply super-resolution algorithms to the joint delay-Doppler estimation problem, the data matrix $\mathbf{X}_{k_{\hat{\theta}}}\in\mathbb{C}^{N\times P}$ in (38) is vectorized as $\mathbf{x}_{k_{\hat{\theta}}}\in\mathbb{C}^{NP\times 1}$, leading to
[…]
@@
where $\mathbf{E}_{n}$ denotes the noise subspace, $\mathbf{a}_{S}\left(\tau,\upsilon\right)=\mathbf{a}_{\upsilon}\left(\upsilon,P_{\mathrm{sub}}\right)\otimes\mathbf{a}_{\tau}\left(\tau,N_{\mathrm{sub}}\right)\in\mathbb{C}^{N_{\mathrm{sub}}P_{\mathrm{sub}}\times 1}$ is the corresponding steering vector, $\mathbf{a}_{\tau}\left(\tau,N_{\mathrm{sub}}\right)$ and $\mathbf{a}_{\upsilon}\left(\upsilon,P_{\mathrm{sub}}\right)$ are the steering vectors in the delay and Doppler domains in (27) and (31), respectively. The peaks of $P_{\mathrm{MUSIC}}\left(\tau,\upsilon\right)$ correspond to the delays and Doppler frequency shifts of the sensing targets.
For the simulation setup in Table V, consider the challenging scenario where the AoAs of the three targets are all $10^{\circ}$, while the other parameters remain unchanged. A comparison of the 2D-periodogram and 2D-MUSIC methods for joint delay-Doppler estimation in this case is shown in Figs. 15(a) and 15(b), where MSSP is perfomed for the 2D-MUSIC algorithm with a window size of $N_{\mathrm{sub}}\times P_{\mathrm{sub}}=\frac{N}{2}\times\frac{P}{2}=64\times 32$. Similar to the 1D-periodogram and 1D-MUSIC, the spectrum of 2D-MUSIC is smoother and exhibits sharper peaks compared to that of the 2D-periodogram. Consequently, 2D-MUSIC offers higher delay and Doppler resolution and can prevent weak targets from being obscured by the sidelobes of strong targets, thereby reducing inter-target interference and achieving better sensing accuracy. However, for 2D-MUSIC, one must correctly choose the dimension of the signal subspace in order to include the impact of the weak targets, and this is a difficult task in general.
@@ IV-D Joint Three-domain Parameter Estimation
In this subsection, we present the joint three-domain parameter estimation method outlined in Fig. 6. Joint AoA-delay-Doppler estimation simultaneously determines the angles, delays, and Doppler frequency shifts of the sensing targets using the signal $\overline{\mathbf{Y}}_{s}$ after element-wise symbol division in (LABEL:d25).
@@
Similar to the 1D and 2D cases, a spectral shift in the angle and Doppler domains of $\mathbf{F}_{\mathrm{3D}}$ to center the zero frequency component within the spectrum is required. Then, from the peak indices $\big(\hat{n}^{k}_{\theta},\hat{n}^{k}_{\tau},\hat{n}^{k}_{\upsilon}\big)$ of $\mathbf{F}_{\mathrm{3D}}$, the angle $\hat{\theta}_{k}$, delay $\hat{\tau}_{k}$, and Doppler frequency $\hat{\upsilon}_{k}$, $\forall k=1,\cdots,K$, of the $K$ sensing targets can be estimated as
@@
TABLE VI: Complexity of 1D parameter estimation algorithms for AoA estimation in MIMO-OFDM ISAC.
@@
$Q{M_{r}^{2}}+M_{r}K\left({M_{r}-K}\right)+\left({2M_{r}\left({M_{r}-K}\right)+M_{r}}\right){N_{\theta}^{\text{s}}}$
@@
$Q{M_{r}^{2}}+M_{r}K\left({M_{r}-K}\right)+{Q^{2}}\left({M_{r}-1}\right)+{K^{3}}$
@@
TABLE VII: The complexity of 1D algorithms along different antenna number for AoA estimation in MIMO-OFDM ISAC.
Algorithms Basic setup and complexity
$N=128,P=64,Q=NP=8192,K=3,N^{s}=N_{\theta}^{\mathrm{IFFT}}=1800$
$M_{r}=16$ $M_{r}=64$ $M_{r}=256$ $M_{r}=512$ $M_{r}=1024$
Theoretical Time(s) Theoretical Time(s) Theoretical Time(s) Theoretical Time(s) Theoretical Time(s)
Periodogram 1.596e8 0.0658 1.596e8 0.0671 1.596e8 0.0785 1.596e8 0.0928 1.596e8 0.1020
MUSIC 2.879e6 0.0020 4.799e7 0.0061 7.874e8 0.0287 3.221e9 0.0946 1.343e10 0.3772
PM-MUSIC 2.876e6 0.0018 4.774e7 0.0056 7.708e8 0.0256 3.088e9 0.0791 1.236e10 0.2790
ROOT-MUSIC / 0.0014 / 0.0110 / 0.3239 / 1.1291 / 4.7785
FFT-MUSIC 2.354e6 0.0015 3.500e7 0.0049 5.586e8 0.0230 2.292e9 0.0704 9.684e9 0.3018
LS-ESPRIT 2.101e6 0.0011 3.382e7 0.0038 5.537e8 0.0239 2.282e9 0.0889 9.664e9 0.4400
TLS-ESPRIT 2.102e6 0.0011 3.382e7 0.0038 5.537e8 0.0247 2.282e9 0.0903 9.664e9 0.4380
PM-ESPRIT 2.098e6 0.0010 3.357e7 0.0033 5.371e8 0.0159 2.148e9 0.0502 8.593e9 0.1900
OMP 7.094e8 0.2543 2.837e9 0.3004 1.135e10 0.5505 2.270e10 0.9933 4.540e10 1.9099
The ESPRIT and OMP algorithms can also be applied to the 2D and 3D parameter estimation problems. Similar to the MUSIC algorithm, spatial smoothing is also required for ESPRIT to achieve decorrelation. The conventional 2D and 3D ESPRIT algorithms require parameter grouping [131] to match the estimation results of the different domains, whereas some modified ESPRIT algorithms [132, 133] can achieve automatic grouping. Furthermore, extensive efforts have been devoted to reducing the complexity of super-resolution methods. For example, the ROOT-MUSIC [134, 135] algorithm replaces the spectrum search in MUSIC with polynomial rooting, the propagator method (PM) [126, 127] uses an alternative approach to achieve the eigenvalue decomposition, and the FFT-MUSIC [128] algorithm accelerates the MUSIC spectrum peak search using FFT operations. In the following, we analyze the complexity and estimation accuracy of the aforementioned algorithms.
IV-E Complexity and Performance Analysis
IV-E1 Complexity Analysis
We examine the theoretical computational complexity of the aforementioned algorithms in terms of their required number of complex-valued multiplications. For the analysis, we assume the complexity associated with multiplying $M\times Q$ and $Q\times N$ matrices is $\mathcal{O}(QMN)$, while the EVD of an $M\times M$ matrix requires $\mathcal{O}(M^{3})$ multiplies. The complexity of various 1D parameter estimation algorithms for AoA estimation is shown in Table VI, and these results also apply to delay and Doppler estimation. In these expressions, $Q$ denotes the number of snapshots, $N^{\mathrm{FFT}}/N^{\mathrm{IFFT}}$ and $N^{\mathrm{s}}$ denote the number of FFT/IFFT points and spectrum search points respectively, which are determined by the estimation interval $\eta$, and $N^{\mathrm{FFT}}=N^{\mathrm{IFFT}}=N^{\mathrm{s}}=\frac{\Delta_{\mathrm{max}}}{\eta}$ where $\Delta_{\mathrm{max}}$ represents the maximum estimation range. The estimation interval $\eta$ affects estimation accuracy, and it should be adjustable in order to achieve specific accuracy requirements. It is difficult to express the theoretical complexity of ROOT-MUSIC due to the difficulty of quantifying the complexity of the polynomial rooting process. For a typical example, we set the number of OFDM subcarriers and symbols as $N=128$ and $P=64$, respectively, the number of targets as $K=3$, the estimation interval as $\eta=0.1^{\circ}$, and the estimation range as $90^{\circ},90^{\circ}]$. Then, the theoretical complexity and code runtime of the various 1D estimation algorithms for different numbers of Sen-RX antennas $M_{r}$ is shown in Table VII, where the code runtime represents the average runtime of implementing the algorithms over 100 iterations. For reference, all algorithms are implemented using MATLAB, and the specific computer specifications are 12-th Generation Intel(R) Core(TM) i7-12700 CPU, 32GB RAM, and 1TB SSD. Given the limitation inherent in the random access machine models for computational complexity [136, 137] and the practical discrepancies in MATLAB operations such as accelerated processing for large-size matrix multiplications, in the following we analyze the complexity of each algorithm based on theoretical complexity, with the time complexity provided only for reference.
TABLE VIII: Complexity of 2D algorithms for joint delay-Doppler estimation and 3D algorithms for joint AoA-delay-Doppler estimation in MIMO-OFDM ISAC.
Algorithms Theoretical complexity
2D-Periodogram $M_{r}(NN_{\upsilon}^{\mathrm{FFT}}\mathrm{log}_{2}N_{\upsilon}^{\mathrm{FFT}}+N_{\upsilon}^{\mathrm{FFT}}N_{\tau}^{\mathrm{IFFT}}\mathrm{log}_{2}N_{\tau}^{\mathrm{IFFT}})$
2D-MUSIC $M_{r}N^{2}P^{2}+{(NP)}^{3}+(2NP\left(NP-K\right)+NP){N_{\tau}^{\text{s}}}{N_{\upsilon}^{\text{s}}}$
2D-ESPRIT $2M_{r}N^{2}P^{2}+{2(NP)}^{3}+K^{2}N\left(P-1\right)+K^{2}P\left(N-1\right)+2K^{3}$
3D-Periodogram $NPN_{\theta}^{\mathrm{IFFT}}\log_{2}N_{\theta}^{\mathrm{IFFT}}+NN_{\theta}^{\mathrm{IFFT}}N_{\upsilon}^{\mathrm{FFT}}\log_{2}N_{\upsilon}^{\mathrm{FFT}}+N_{\theta}^{\mathrm{IFFT}}N_{\upsilon}^{\mathrm{FFT}}N_{\tau}^{\mathrm{IFFT}}\log_{2}N_{\tau}^{\mathrm{IFFT}}$
3D-MUSIC ${(M_{r}NP)}^{2}+{(M_{r}NP)}^{3}+(2M_{r}NP\left(M_{r}NP-K\right)+M_{r}NP){N_{\theta}^{\text{s}}}{N_{\tau}^{\text{s}}}{N_{\upsilon}^{\text{s}}}$
(a) AoA estimation RMSE.
(b) Delay estimation RMSE.
(c) Doppler estimation RMSE.